a) Find the density of air (in kg/m3) at a pressure of 1.00 atm and a temperature of 35.0°C, assuming that air is 78% N2, 21% O2, and 1% Ar. kg/m3

b) Find the density of the atmosphere on Venus (in kg/m3), assuming that it's 96% CO2 and 4% N2, with a temperature of 724 K and a pressure of 92.0 atm.
Answer #1

a]Molar mass of air = [0.78*28+0.21*32+0.01*40] =28.96 g

By using ideal gas equation ,

V = nRT/P = 1*8.314*(273+35)/(1*101325) = 0.02527 m^3

density = mass/ volume = 28.96e-3/0.02527

= 1.146 kg/m^3

b] Molar mass of air = [0.96*44+0.04*28] = 43.36 g

By using ideal gas equation ,

V = nRT/P = 1*8.314*724/(92*101325) = 0.0006457 m^3

density = mass/ volume = 43.36e-3/0.0006457

= 67.15 kg/m^3